# Less but Better: Enabling Generalized Zero-shot Learning Towards Unseen Domains by Intrinsic Learning from Redundant LLM Semantics

### v4

@@ IV-B Experimental details
IV-C Results on Office-home and Mini-DomainNet
We have adopted six comparative methods to compare with the proposed MDASR. Specifically, these six methods can be divided into CDZSL methods (including CuMix[21] and its two variants (img-only and two-level), S-AGG [35], and SEIC [23]) and generative method modified to apply in various domains (f-CLSWGAN [13]). CDZSL methods encompass the current approaches that aim to identify unseen classes in unseen domains. However, due to the influence of the domain shift problem, they are not very effective at recognizing both seen and unseen classes simultaneously. Comparatively, generative methods can better overcome the impact of domain shift problem, but it is not designed for CDGZSL. These comparative methods utilize semantics in two ways: when semantics exist in the form of an attribute matrix, they can be used directly; when semantics are in text form, they can first be converted into sentence vectors using sentence-Bert, and then utilized. Since our proposed method is a generative approach with meta semantic refinement, we also employed PCA[36] as another semantic refinement method for comparison.
@@ IV-C Results on Office-home and Mini-DomainNet
For the Office-home dataset, we utilized two types of semantic sources: one acquired from querying LLMs and the other manually annotated. The results for both types of semantics are displayed in Table II. It is evident that the proposed MDASR outperforms other methods with both LLM-provided semantics and manually annotated semantics, with the former even surpassing the performance of the latter. We believe this is because the Office-home dataset consists of everyday items, which LLM is already capable of describing comprehensively, whereas manually provided semantics are limited by the annotators’ energy and ability to provide a thorough description. However, expert intervention may still be necessary in some non-daily scenarios like industry and medicine. Observing the comparative methods, the existing four CDZSL methods, CuMix and its variants (”img-only” and ”two-level” respectively refer to performing mixup only in the image space and performing mixup in both the image and feature spaces), S-AGG, and SEIC, all have a lower accuracy for unseen classes due to domain shift problem. In comparison, generative methods can somewhat overcome this issue. It can be observed that the modified f-CLSWGAN is the current state-of-the-art. Meanwhile, the PCA semantic refinement used for comparison has negative effects on generative methods most time. The performance can often tend towards degradation. This is because the optimization of PCA does not consider the effect of semantic redundancy, and does not take measurement to preserve effective semantics. The proposed MDASR, in contrast, enhances performance over f-CLSWGAN across different semantics and domains.
TABLE III: The CDGZSL results on Mini-DomainNet dataset
{tblr}
cell11 = c=2,r=2, cell13 = c=3, cell16 = c=3, cell19 = c=3, cell31 = r=3, cell61 = c=2, cell71 = c=2, cell81 = r=2, cell111 = c=2, hline1,3,11 = -, hline2 = 3-11, hline4-5,9-10 = 2, hline6-8,10-11 = 1-2, Methods & Clipart Painting Sketch
S U H S U H S U H
CuMix img-only 63.98 2.67 5.12 63.69 2.64 5.06 59.61 1.86 3.60
two-level 64.54 3.32 6.31 64.20 2.58 4.96 59.75 3.42 6.46
original 64.20 2.49 4.79 64.27 3.10 5.91 54.70 5.07 9.27
S-AGG 63.98 2.89 5.53 62.65 0.90 1.77 56.96 1.91 3.69
SEIC 64.24 2.01 3.89 63.46 1.09 2.14 58.45 1.08 2.12
f-CLSWGAN original 61.38 24.53 35.05 58.23 23.31 33.29 56.60 22.03 31.71
PCA 64.51 22.27 33.10 66.73 11.10 19.03 58.79 8.62 15.03
MDASR 63.69 29.43 40.25 61.02 26.16 36.62 55.16 28.25 37.36
@@
For the Mini-DomainNet dataset, due to the larger number of classes, we only demonstrate the results using semantics automatically acquired from LLMs, which are presented in Table III. Given that Mini-DomainNet has a greater number of unseen classes, the difficulty of CDGZSL is also higher. The proposed MDASR achieves an H-accuracy ranging from 36.62% to 40.25% in different unseen domains, which remains the highest among all the methods compared. The results of the other methods are similar to those on Office-Home but with a general decline in H-accuracy. The domain shift problem effect is still severe with CDZSL methods. Under the CDZSL setting, their accuracy for unseen classes may reach 8.50%-25.90% [23], but under the CDGZSL setting, it’s only about 2%-5%. f-CLSWGAN still maintains a certain capacity for alleviating domain shift problem. If our MDASR is leveraged, the H-accuracy is further improved compared to f-CLSWGAN; however, if PCA is used for semantic refinement, since its objective function is to maximize variance, which is unrelated to the intrinsic semantics, it usually brings about a negative effect. Overall, in the two datasets, our proposed MDASR achieved the highest H-accuracy, presenting the capacity to recognize seen and unseen classes in unseen domains.
@@
Additionally, to further illustrate the effectiveness of the proposed MDASR in eliminating redundant semantics, we performed a two-dimensional visualization of the original semantic embeddings and the refined semantic embeddings, as shown in Fig. 7. Since the Mini-domainnet category count exceeds 100, considering spatial constraints and presentation effectiveness, we randomly selected 17 categories for display. From the figure, it can be seen that the refined semantics are more closely aligned with intrinsic visual features rather than other redundant information. For example, initially, in the original semantics, watermelon and broccoli are very close together but far from potato, which may suggest a large proportion of color in the original semantics since the first two are green. However, after refinement, the distance between watermelon and potato has decreased, indicating that the new semantics are no longer influenced by color, which can easily change with domain, but instead refer to the more difficult to alter ’shape,’ as both are elliptically shaped. Furthermore, in the original semantics, vase and candle nearly overlap, which is inconsistent with visual features; the refined semantics also separate the two. Lastly, in the original semantics, butterfly, mosquito, and spider are far apart, but the refined semantics bring them closer together, which is also consistent with visual features. Overall, our method is effective in eliminating redundant components in semantics under the supervision of the common feature space.
TABLE IV: The ablation experiment results on Mini-DomainNet dataset
{tblr}
cell11 = c=3, cell14 = c=3, cell17 = c=3, cell110 = c=3, hline1,3,7 = -, hline2 = 2-13, & Loss Clipart Painting Sketch
$\mathcal{L}_{s}$ $\mathcal{L}_{meta}$ S U H S U H S U H
Case (a): ✘ ✘ 61.38 24.53 35.05 58.23 23.31 33.29 56.60 22.03 31.71
Case (b): ✘ 61.34 29.32 39.67 58.29 20.23 30.03 57.32 26.02 35.79
Case (c): ✘ 63.73 26.15 37.08 59.98 23.82 34.10 55.41 26.25 35.62
Case (d): 63.69 29.43 40.25 61.02 26.16 36.62 55.16 28.25 37.36
IV-D Ablation Study
@@
We conducted ablation experiments on the Mini-DomainNet dataset to prove the effectiveness of each component in our proposed method. The relevant results are presented in Table IV. There are a total of four sets of results, corresponding to whether similarity alignment loss $\mathcal{L}_{s}$ and meta-learning loss $\mathcal{L}_{meta}$ are used or not. Case (a) represents a generative method without any semantic refinement in the aligned space, and its results achieve 33.35% on average. Next, we verified the two parts of the proposed meta semantic refinement loss functions: the similarity matrix alignment $\mathcal{L}_{s}$ and the meta-learning loss $\mathcal{L}_{meta}$. It can be observed that using only the similarity matrix alignment, as in case (b), some results improve while others worsen, suggesting that this method might lose certain important semantics. Solely utilizing meta-learning does not lead to degradation but offers limited improvement, as shown in case (c). Only by using both, as in case (d), do we achieve the best semantic refinement results, which the average accuracy is 38.07%.
V Conclusion
@@
In this paper, we have introduced the task of cross-domain generalized zero-shot learning towards unseen domains for the first time and developed the MDASR approach which addresses the information asymmetry caused by the redundancy of LLM-based semantics. We construct a common feature space of seen domains, enabling the discriminative relations learned from the seen domains to generalize to unseen domains. Under the guidance of the common feature space, by designing ISA which aligns intra-class relationship, we have achieved the removal of non-intrinsic semantics. Meanwhile, through UMG that simulates the synthesis of unseen features, we retain intrinsic semantics that could connect the seen and unseen classes. Building on this, our proposed method can generate features of unseen classes by leveraging intrinsic semantics shared across domains, alleviating domain shift problem and recognizing both seen and unseen classes in unseen domains. We have validated the CDGZSL capabilities of our proposed method on two open-source datasets. On the Office-Home dataset, the average H-accuracy of CDGZSL reached 47.11%, and on the more challenging Mini-DomainNet dataset, the average H-accuracy of CDGZSL reached 38.07%. We have also made the corresponding semantics available to the public, providing a benchmark for future researchers.

### v5

@@ 4.2 Experimental details
4.3 Results on Office-home and Mini-DomainNet
We have adopted six comparative methods to compare with the proposed MDASR. Specifically, these methods can be divided into two groups: CDZSL methods and recent generative methods for zero-shot learning. The CDZSL methods include CuMix(Mancini et al., 2020) (and its two variants, “img-only” and “two-level”, which respectively refer to performing mixup only in the image space and performing mixup in both the image and feature spaces), S-AGG(Chandhok et al., 2021), and SEIC(Mondal and Biswas, 2022). These methods aim to identify unseen classes in unseen domains but are significantly affected by the domain shift problem, leading to poor performance on unseen classes under CDGZSL settings. The generative methods we adopted for comparison include two recent state-of-the-art zero-shot learning methods: ZeroDiff(Ye et al., 2025) and VADS(Hou et al., 2024), which we adapted to CDGZSL. These generative methods have demonstrated strong capabilities in alleviating the domain shift problem. All comparative methods utilize semantics as follows: when semantics are provided in the form of attribute matrices, they are used directly; when semantics are textual, they are first transformed into semantic embeddings through sentence-Bert and subsequently utilized for prediction.
@@ 4.3 Results on Office-home and Mini-DomainNet
For the Office-home dataset, we utilized two types of semantic sources: one acquired from querying LLMs and the other manually annotated. The results for both types of semantics are displayed in Table 3. It is evident that the proposed MDASR outperforms other methods with both LLM-provided semantics and manually annotated semantics, with the former even surpassing the performance of the latter. We believe this is because the Office-home dataset consists of everyday items, which LLM is already capable of describing comprehensively, whereas manually provided semantics are limited by the annotators’ effort and ability to provide a thorough description. However, expert intervention may still be necessary in some non-daily scenarios like industry and medicine. Observing the comparative methods, the existing four CDZSL methods, CuMix and its variants (“img-only” and “two-level” respectively refer to performing mixup only in the image space and performing mixup in both the image and feature spaces), S-AGG, and SEIC, all have a lower accuracy for unseen classes due to domain shift problem.
Among generative methods, ZeroDiff and VADS represent the recent state-of-the-art approaches that have shown promising performance in the zero-shot learning scenario. However, when adapted to the CDGZSL setting, their performances vary significantly across different domains and semantic types. Specifically, ZeroDiff generally achieves a relatively balanced performance across both semantics, while VADS exhibits notable variance in its ability to handle different semantics and domains. Nevertheless, both methods still show limited accuracy for unseen classes in unseen domains, indicating that directly applying these generative ZSL methods into CDGZSL tasks remains challenging. In contrast, our proposed MDASR consistently surpasses ZeroDiff and VADS by a considerable margin, demonstrating the effectiveness of the proposed meta semantic refinement strategy in handling semantic redundancy and domain shift simultaneously.
Table 4: The CDGZSL results on Mini-DomainNet dataset
Methods Clipart Painting Sketch S U H S U H S U H CuMix img-only 63.98 2.67 5.12 63.69 2.64 5.06 59.61 1.86 3.60 two-level 64.54 3.32 6.31 64.20 2.58 4.96 59.75 3.42 6.46 original 64.20 2.49 4.79 64.27 3.10 5.91 54.70 5.07 9.27 S-AGG 63.98 2.89 5.53 62.65 0.90 1.77 56.96 1.91 3.69 SEIC 64.24 2.01 3.89 63.46 1.09 2.14 58.45 1.08 2.12 ZeroDiff 62.32 3.87 7.29 60.61 5.01 9.26 57.27 5.07 9.31 VADS 63.52 4.65 8.66 56.28 6.21 11.19 56.59 3.94 7.37 MDASR 63.69 29.43 40.25 61.02 26.16 36.62 55.16 28.25 37.36
    The bold values represent the highest H-accuracy in each part.
@@
For the Mini-DomainNet dataset, due to the larger number of classes, we only demonstrate the results using semantics automatically acquired from LLMs, which are presented in Table 4. Given that Mini-DomainNet has a greater number of unseen classes, the difficulty of CDGZSL is also higher. The proposed MDASR achieves an H-accuracy ranging from 36.62% to 40.25% in different unseen domains, which remains the highest among all the methods compared. The results of the other methods are similar to those on Office-Home but with a general decline in H-accuracy. The domain shift problem effect is still severe in CDZSL methods. Under the CDZSL setting, their accuracy for unseen classes may reach 8.50%-25.90% (Mondal and Biswas, 2022), but under the CDGZSL setting, it is only about 2%-5%. Among the generative zero-shot learning methods, ZeroDiff and VADS exhibit moderate performance in alleviating the domain shift problem on Mini-DomainNet. However, due to the complexity and large category number of Mini-DomainNet, the improvement brought by these methods is still limited under the CDGZSL setting. In contrast, the proposed MDASR significantly outperforms ZeroDiff and VADS, showcasing strong generalization across various domains and semantics. This result further validates the robustness and effectiveness of our meta semantic refinement method in learning intrinsic semantic representations.
@@
Additionally, to further illustrate the effectiveness of the proposed MDASR in eliminating redundant semantics, we performed a two-dimensional visualization of the original semantic embeddings and the refined semantic embeddings, as shown in Fig. 7. Since the Mini-domainnet category count exceeds 100, considering spatial constraints and presentation effectiveness, we randomly selected 17 categories for display. From the figure, it can be seen that the refined semantics are more closely aligned with intrinsic visual features rather than other redundant information. For example, initially, in the original semantics, watermelon and broccoli are very close to each other but far from potato, which may suggest a large proportion of color in the original semantics since the first two are green. However, after refinement, the distance between watermelon and potato has decreased, indicating that the new semantics are no longer influenced by color, which can easily change with domain, but instead refer to ‘shape’ that is more diffcult to alter as both are elliptically shaped. Furthermore, in the original semantics, vase and candle nearly overlap, which is inconsistent with visual features; the refined semantics also separate the two. Lastly, in the original semantics, butterfly, mosquito, and spider are far apart, but the refined semantics bring them closer to each other, which is also consistent with visual features. Overall, our method is effective in eliminating redundant components in semantics under the supervision of the common feature space.
4.4 Ablation Study
Figure 8: The ablation experiment results on Mini-DomainNet dataset
@@ 4.4 Ablation Study
We conducted ablation experiments on the Mini-DomainNet dataset to prove the effectiveness of each component in our proposed method. The relevant results are presented in Fig. 8. We conducted a total of four groups of experiments: Case (a): without ISA and UMG, Case (b): with ISA, without UMG, Case (c): without ISA, with UMG, Case (d): with both ISA and UMG. Specifically, case (a) represents a generative method without any semantic refinement in the aligned space, and it results in 33.35% accuracy on average. Next, we verified the two parts of the proposed meta semantic refinement loss functions: the similarity matrix alignment $\mathcal{L}_{s}$ and the meta-learning loss $\mathcal{L}_{meta}$. It can be observed that using only the similarity matrix alignment, as in case (b), some results improve while others worsen, suggesting that this method might lose certain important semantics. Solely utilizing meta-learning does not lead to degradation but offers limited improvement, as shown in case (c). Only by using both, as in case (d), do we achieve the best semantic refinement results, with the average accuracy of 38.07%.
5 Conclusion
@@
In this paper, we have introduced the task of cross-domain generalized zero-shot learning towards unseen domains for the first time and developed the MDASR approach which addresses the information asymmetry caused by the redundancy of LLM-based semantics. We construct a common feature space of seen domains, enabling the discriminative relations learned from the seen domains to generalize to unseen domains. Under the guidance of the common feature space, by designing ISA which aligns intra-class relationship, we have achieved the removal of non-intrinsic semantics. Meanwhile, through UMG that simulates the synthesis of unseen features, we retain intrinsic semantics that could connect the seen and unseen classes. Building on this, our proposed method can generate features of unseen classes by leveraging intrinsic semantics shared across domains, alleviating domain shift problem and recognizing both seen and unseen classes in unseen domains. We have validated the CDGZSL capabilities of the proposed method on two open-source datasets. On the Office-Home dataset, the average H-accuracy of CDGZSL reached 47.11%, and on the more challenging Mini-DomainNet dataset, the average H-accuracy of CDGZSL reached 38.07%. We have also made the corresponding semantics available to the public, providing a benchmark for future studies.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (No. 62450020), and in part by the National Science Fund for Distinguished Young Scholars (No. 62125306), and in part by Zhejiang Key Research and Development Project (2024C01163), and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2024B19).